# MACE-Dance: Motion-Appearance Cascaded Experts for Music-Driven Dance Video Generation

### v1

@@ 4.1 Dataset
Table 3: Quantitative comparison with SOTAs on the MA-Data dataset in Pose-Driven Image Animation task.
@@
Animate-Anyone [12] [CVPR’24] 515.26 0.648 0.091 19.65
Magic-Animate [43] [CVPR’24] 1032.06 0.311 0.207 14.00
Wan-Animate [5] [Arxiv’25] 332.82 0.707 0.078 21.11
@@
Figure 3: Qualitative comparison with SOTAs across reference image domains (real-person vs. anime-character) and music genres (Eastern Folk vs. Popping) in the music-driven dance video generation task.
4.2 Evaluation
The key challenges of music-driven dance video generation are: (1) generating dance motions that are kinematically plausible while artistically expressive; and (2) achieving high-fidelity visual appearance with strong spatiotemporal consistency. Inspired by this, we introduce a motion–appearance evaluation protocol. (1) Motion dimension. We first extract 2D keypoint sequences using ViTPose [41] from the dance videos and evaluate from a Human-Kinematics perspective. To evaluate the fidelity and diversity, we report FID and DIV across two feature spaces [16, 30]: (1) kinetic (k), capturing motion dynamics, and (2) geometric (g), encoding spatial joint relations. To measure music–motion synchronization, we utilize the Beat Alignment Score (BAS) [16, 19]. (2) Appearance dimension. We adopt a set of metrics from VBench [14], specifically choosing those most critical for the nuances of dance videos. Our evaluation includes imaging quality (IQ), aesthetic quality (AQ), subject consistency (SC), background consistency (BC), motion smoothness (MS), temporal flickering (TF). Finally, we collect an additional 500 5-second clips to construct the test set, with high engagement on TikTok and across multiple dance genres.
4.3 Comparison
@@
Music-Driven Dance Video Generation. As there is currently no open-source implementation for Music-Driven Dance Video Generation, we compare MACE-Dance against two baseline families on the MA-Data dataset: (1) 3D dance generation methods pipelined with Wan-Animate, including EDGE [37], Lodge [19], and MEGA [47]; (2) talking-head generation methods adapted by replacing face masks with full-body masks followed by fine-tuning on the MA-Data dataset, exemplified by Hallo2 [6]. As shown in Tab. 1, our proposed MACE-Dance demonstrates state-of-the-art (SOTA) performance in both appearance and motion quality. Specifically, for the motion aspect, MACE-Dance achieves the best results across all metrics ($FIK_{k}=16.46,FID_{g}=0.28,DIV_{k}=9.74,DIV_{g}=6.34,BAS=0.523$); for the appearance aspect, it also attains best performance on most metrics, with scores of IQ=65.35,AQ=51.79,SC=93.97,BC=94.57,MS=98.46,IQ=65.35,AQ=51.79,SC=93.97,BC=94.57,MS=98.46, and $TF=97.10$. By effectively decoupling the task into 3D Dance Generation and Pose-Driven Image Animation, and leveraging the strong performance of the Motion Expert and Appearance Expert on their respective sub-tasks, MACE-Dance delivers exceptional generation quality. Note: User Study can be found in the supplementary material.
@@
Music-Driven 3D Dance Generation. Music-Driven 3D Dance Generation is a canonical task and determines the motion quality of MACE-Dance. We compare the Motion Expert against FACT [16], MNET [15], Bailando [30], EDGE [37], Lodge [19], and MEGA [47] on the FineDance dataset with metrics following  [19]. As shown in Tab. 2, the Motion Expert attains overall state-of-the-art (SOTA) performance. Specifically, it achieves the best $FID_{k}=17.83$ and a competitive $FID_{g}=25.09$, indicating high fidelity; the best $DIV_{k}=10.30$ and $DIV_{g}=8.09$, indicating strong diversity; a competitive FSR, supporting physical plausibility; the best $BAS=0.229$, demonstrating superior audio-motion synchronization; and a substantially higher $FPS=770$, evidencing excellent generation efficiency. These advances primarily stem from: (1) adopting a Diffusion Model with a BiMamba-Transformer hybrid architecture, enabling high-quality long motion sequences in a non-autoregressive manner; and (2) a Guidance-Free Training (GFT) strategy that improves generation quality without requiring dual-pass inference (conditioned + unconditioned).
@@
Pose-Driven Image Animation. Pose-Driven Image Animation is likewise a canonical task and governs the appearance quality of MACE-Dance. We compare the Appearance Expert against Animate-Anyone [12], Magic-Animate [43], and Wan-Animate [5] on the MA-Data dataset with metrics following  [5]. As shown in Tab. 3, the Appearance Expert achieves state-of-the-art (SOTA) performance, on all metrics FVD=274.94,SSIM=0.739,LPIPS=0.066,PSNR=22.40FVD=274.94,SSIM=0.739,LPIPS=0.066,PSNR=22.40. Its strong video synthesis quality primarily benefits from Wan-Animate (Baseline)’s powerful cross-modal understanding and our proposed Kinematic-Aesthetic decoupled fine-tuning strategy. Note: Qualitative analysis can be found in the supplementary material.
4.4 Qualitative Analysis
@@
We also present a qualitative comparison against other methods across reference-image domains (real person vs. anime character) and music genres (elegant and rhythmically rich Eastern Folk vs. powerful and funk-inspired Popping) for the music-driven dance video generation task. As shown in Fig.  3, Hallo2 exhibits significant blurring in human details and noticeable artifacts in the background ; EDGE frequently produces abnormal movements that violate physical plausibility; Lodge often exhibits abrupt motion discontinuities, leading to poor visual appearance; and MEGA tends to generate overly simple and repetitive motions that lack artistic expressiveness. In contrast, videos generated by our method (MACE-Dance) not only present kinematically plausible and artistically expressive motion, but also maintain spatiotemporally coherent appearance. In summary, MACE-Dance demonstrates superior performance in qualitative evaluations.
Figure 4: Ablation for the model architecture of Motion Expert.
4.5 Ablation
@@
Figure 5: Ablation for the Appearance Expert.
@@
Appearance Expert Since the two-stage fine-tuning strategy is central to our Appearance Expert, we evaluate the contribution of each stage via ablation, as summarized in Tab.  3 and Fig. 5. (1) Kinematic Stage. We fine-tune the Body Adapter while freezing all other components. Removing this stage leads to a modest decline across all metrics quantitatively and noticeable kinematic errors and motion blur qualitatively, indicating its effectiveness for ensuring human kinematic plausibility in video. (2) Aesthetic Stage. We fine-tune LoRA parameters in each DiT block. Omitting this stage causes a substantial degradation across all metrics quantitatively and obvious ghosting artifact qualitatively, underscoring its critical role to preserve video aesthetic. Finally, Appearance Expert also demonstrates the superior performance over the Wan-Animate (Baseline), which validates the overall effectiveness of the proposed Kinematic-Aesthetic fine-tuning strategy.
5 Conclusion
@@
In conclusion, we present MACE-Dance, a music-driven dance video generation framework with cascaded Mixture-of-Experts (MoE). The Motion Expert enforces kinematic plausibility and artistic expressiveness, while the Appearance Expert preserves visual identity with spatiotemporal coherence. Specifically, the Motion Expert adopts Diffusion Model with a BiMamba–Transformer hybrid backbone and Guidance-Free Training strategy, while the Appearance Expert adopts a decoupled Kinematic–Aesthetic fine-tuning strategy. To better benchmark this task, we curate a large-scale and diverse dataset, and desigm a motion–appearance evaluation protocol.Extensive experiments demonstrate the superiority of MACE-Dance and of its Motion and Appearance Experts. For future work, we plan to extend MACE-Dance with textual descriptions to enable more interactive and flexible dance generation, as text serves as the most flexible and intuitive form for specifying user needs.

### v3

@@ 4.2. Evaluation
Table 3. Quantitative comparison with SOTAs on the MA-Data dataset in Pose-Driven Image Animation task.
@@
Animate-Anyone (Hu, 2024) [CVPR’24] 515.26 0.648 0.091 19.65
Magic-Animate (Xu et al., 2024a) [CVPR’24] 1032.06 0.311 0.207 14.00
Wan-Animate (Cheng et al., 2025) [Arxiv’25] 332.82 0.707 0.078 21.11
@@
Figure 4. MACE-Dance generates high-quality dance videos across diverse dance genres.
4.3. Comparison
4.3.1. Music-Driven Dance Video Generation.
@@ 4.3. Comparison
As there is currently no open-source implementation for Music-Driven Dance Video Generation, we compare MACE-Dance against two baseline families on the MA-Data dataset: (1) 3D dance generation methods pipelined with Wan-Animate, including EDGE (Tseng et al., 2023), Lodge (Li et al., 2024c), and MEGA (Yang et al., 2025b); (2) General human-motion video generation methods. We perform inference using the pretrained EchoMimic-V3 (Meng et al., 2025) and WAN-S2V (Gao et al., 2025) models, which supports general human motion generation. In addition, we adapt the Hallo2 (Cui et al., 2024) by replacing facial masks with full-body masks, and then fine-tune it on the MA-Data dataset. As shown in Tab.  1, our proposed MACE-Dance demonstrates state-of-the-art (SOTA) performance in both appearance and motion quality. Specifically, for the motion aspect, MACE-Dance achieves the best results across all metrics ($FID_{k}=16.46,FID_{g}=0.28,DIV_{k}=9.74,DIV_{g}=6.34,BAS=0.523$); for the appearance aspect, it also attains best performance on most metrics, with scores of IQ=65.35,AQ=51.79,SC=93.97,BC=94.57,MS=98.46,IQ=65.35,AQ=51.79,SC=93.97,BC=94.57,MS=98.46, and $TF=97.10$. By effectively decoupling the task into 3D Dance Generation and Pose-Driven Image Animation, and leveraging the strong performance of the Motion Expert and Appearance Expert on their respective sub-tasks, MACE-Dance delivers exceptional generation quality. Note: User Study can be found in the supplementary material Sec. 2.
4.3.2. Music-Driven 3D Dance Generation.
@@
Music-Driven 3D Dance Generation is a canonical task and determines the motion quality of MACE-Dance. We compare the Motion Expert against FACT (Li et al., 2021), MNET (Kim et al., 2022), Bailando (Siyao et al., 2022), EDGE (Tseng et al., 2023), Lodge (Li et al., 2024c), and MEGA (Yang et al., 2025b) on the FineDance dataset with metrics following  (Li et al., 2024c). As shown in Tab.  2, the Motion Expert attains overall state-of-the-art (SOTA) performance. Specifically, it achieves the best $FID_{k}=17.83$ and a competitive $FID_{g}=25.09$, indicating high fidelity; the best $DIV_{k}=10.30$ and $DIV_{g}=8.09$, indicating strong diversity; a competitive FSR, supporting physical plausibility; the best $BAS=0.229$, demonstrating superior audio-motion synchronization; and a substantially higher $FPS=770$, evidencing excellent generation efficiency. These advances primarily stem from: (1) adopting a Diffusion Model with a BiMamba-Transformer hybrid architecture, enabling high-quality long motion sequences in a non-autoregressive manner; and (2) a Guidance-Free Training (GFT) strategy that improves generation quality without requiring dual-pass inference (conditioned + unconditioned). Note: Qualitative Comparison can be found in the supplementary material Sec. 3.2.
4.3.3. Pose-Driven Image Animation.
@@
Pose-Driven Image Animation is likewise a canonical task and governs the appearance quality of MACE-Dance. We compare the Appearance Expert against Animate-Anyone (Hu, 2024), Magic-Animate (Xu et al., 2024a), and Wan-Animate (Cheng et al., 2025) on the MA-Data dataset with metrics following  (Cheng et al., 2025). The Appearance Expert achieves state-of-the-art (SOTA) performance on all metrics (Tab. 3, $FVD=274.94$, $SSIM=0.739$, $LPIPS=0.066$, $PSNR=22.40$). Its strong video synthesis quality primarily benefits from Wan-Animate (Baseline)’s powerful cross-modal understanding and the Kinematic-Aesthetic decoupled fine-tuning strategy. Note: Qualitative Comparison can be found in the supplementary material Sec. 3.3.
Figure 5. Ablation for the model architecture of Motion Expert.
Figure 6. MACE-Dance produces coherent long-sequence dance videos.
Figure 7. Ablation for the Appearance Expert.
4.4. Qualitative Analysis
4.4.1. Effect Comparison.
@@ 4.4. Qualitative Analysis
We also present a qualitative comparison against other methods across reference-image domains (real person vs. anime character) and music genres (elegant and rhythmically rich Eastern Folk vs. powerful and funk-inspired Popping) for the music-driven dance video generation task. As shown in Fig.  3, Hallo2 exhibits significant blurring in human details and noticeable artifacts; EDGE often shows abrupt motion discontinuities; Lodge frequently produces abnormal movements that violate physical plausibility; MEGA, WAN-S2V, and Echomimic-V3 often produce overly simple and repetitive motions, limiting expressiveness. In contrast, videos generated by MACE-Dance not only present kinematically plausible and artistically expressive human motion, but also maintain spatiotemporally coherent visual appearance.
4.4.2. Cross-Genre Generation.
Moreover, MACE-Dance generalizes effectively across dance genres as shown in Fig.  4, producing distinct genre-specific motion signatures, including (1) Uyghur dance exhibits light, continuous upper-body rotations with expressive arm trajectories; (2) Dunhuang motion features stable lower-body stances and elegant, circular arm patterns; (3) Dai style emphasizes soft, flowing wrist and elbow movements; (4) K-Pop example demonstrates crisp transitions, symmetrical poses, and rhythm-driven gestures; (5) Popping is characterized by sharp isolations and staccato movements, reflecting its percussive movement vocabulary.
4.4.3. Long-Sequence Generation.
Additionally, a complete music track typically lasts 30 seconds to 5 minutes, making long-sequence generation crucial for practical dance video synthesis (Feng et al., 2025). To mitigate motion drift or visual degradation in long-sequence generation, MACE-Dance incorporates dedicated designs in both stages: (1) a BiMamba–Transformer hybrid in the Motion Expert for drift-free long motion synthesis, and (2) pose-driven relay rendering with identity anchoring in the Appearance Expert. As shown in Fig.  14, MACE-Dance produces coherent long-sequence dance video.
4.5. Ablation Study
4.5.1. Motion Expert
@@
4.5.2. Appearance Expert
@@
Since the two-stage fine-tuning strategy is central to our Appearance Expert, we evaluate the contribution of each stage via ablation, as summarized in Tab.  3 and Fig.  7. (1) Kinematic Stage. We fine-tune the Body Adapter while freezing all other components. Removing this stage leads to a modest decline across all metrics quantitatively and noticeable kinematic errors and motion blur qualitatively, indicating its effectiveness for ensuring human kinematic plausibility in video. (2) Aesthetic Stage. We fine-tune LoRA parameters in each DiT block. Omitting this stage causes a substantial degradation across all metrics quantitatively and obvious ghosting artifact qualitatively, underscoring its critical role to preserve video aesthetic. Finally, Appearance Expert also demonstrates the superior performance over the Wan-Animate (Baseline), which validates the overall effectiveness of the proposed Kinematic-Aesthetic fine-tuning strategy.
4.5.3. Motion Representation (2D vs. 3D)
Most pose-driven image animation methods rely on 2D keypoints, which naturally motivates using 2D poses as the intermediate motion representation. In contrast, MACE-Dance adopts 3D motion as its intermediate representation. To validate this design, we compare 2D and 3D representations at both the motion-generation level and the final video-generation level. Specifically, for the motion level, we train the same Motion Expert on FineDance with either 2D or 3D motion targets. For the video level, we further render final videos on MA-Data using either 2D pose sequences directly or 3D motion projected to 2D via our projector. As shown in Tab. 4, the 3D-based representation consistently outperforms the 2D-based one across both settings. On FineDance, 3D achieves better fidelity, diversity, and synchronization, indicating that it provides more stable and physically consistent supervision for music-to-motion learning. On MA-Data, 3D further yields clearly better subject consistency, visual fidelity, and beat alignment in the final rendered videos, showing that the advantages of 3D are preserved after the downstream animation stage. These results demonstrate that 3D motion serves as a more reliable intermediate interface than 2D pose for both controllable dance generation and high-quality video synthesis.
Table 4. Comparison of 2D and 3D motion representations.
Kinematic Appearance
[…]
2D 22.8 8.6 6.12 5.24 0.527 51.86 91.84 23.73 0.496
3D 19.5 4.1 8.87 5.92 0.543 51.79 93.97 16.46 0.523
Table 5. Cross-composition analysis of the two experts on MA-Data.
[…]
w/o.ME 50.21 92.10 20.84 0.499
w/o.AE 50.36 91.42 17.92 0.519
Ours 51.79 93.97 16.46 0.523
4.5.4. Role of Each Expert
To further analyze the role of each expert in MACE-Dance, we conduct an additional cross-composition study on MA-Data by replacing one expert at a time with its corresponding baseline counterpart. Specifically, w/o.ME denotes the variant that uses the baseline Motion Expert (EDGE (Tseng et al., 2023)) together with our Appearance Expert, while w/o.AE denotes the variant that uses our Motion Expert together with the baseline Appearance Expert (WAN-Animate (Cheng et al., 2025)). As shown in Tab. 5, the full MACE-Dance consistently achieves the best performance across all evaluated metrics, confirming that both experts contribute positively to the final music-driven dance video generation quality. More specifically, replacing our Appearance Expert with the baseline model leads to clear degradation in appearance-related metrics such as AQ and SC, while replacing our Motion Expert results in a more noticeable drop in the motion-related metric BAS. These observations suggest that the two experts play complementary roles: the Motion Expert mainly strengthens music-motion alignment and body dynamics, whereas the Appearance Expert further improves visual quality, temporal coherence, and identity consistency in the rendered videos.
4.6. Comparison with Video Foundation Models
We further compare MACE-Dance with general-purpose video foundation models, including CogVideoX1.5-5B (Yang et al., 2024d) and WAN2.2-5B (Wan et al., 2025), to examine how a structured motion-to-appearance pipeline performs against recent large-scale video generation models. Although these models demonstrate strong generation ability in broad video domains, they are not specifically designed for music-driven dance video generation, where accurate modeling of beat, rhythm, and body-motion coherence is particularly important. As shown in Table 6, MACE-Dance achieves the best overall performance on SC, FID, and BAS, indicating stronger music-motion alignment, better visual quality, and more consistent identity preservation. Although WAN2.2-5B attains a slightly higher AQ score, it underperforms our method on the other three metrics. Qualitatively, CogVideoX1.5-5B tends to produce weaker and slower dance motions with noticeable blur, while WAN2.2-5B generates larger motion amplitudes but often suffers from temporal identity inconsistency, as shown in Fig. 8. Overall, these results support the effectiveness of explicitly decomposing music-driven dance video generation into a Motion Expert and an Appearance Expert.
Table 6. Comparison with general-purpose video foundation models.
[…]
CogVideoX1.5-5B 50.38 89.92 22.47 0.477
WAN2.2-5B 53.22 90.77 17.53 0.452
Ours 51.79 93.97 16.46 0.523
Figure 8. Comparison with general-purpose video foundation models.
5. Conclusion
@@
In conclusion, we present MACE-Dance, a music-driven dance video generation framework with cascaded Mixture-of-Experts (MoE). The Motion Expert enforces kinematic plausibility and artistic expressiveness, while the Appearance Expert preserves visual identity with spatiotemporal coherence. Specifically, the Motion Expert adopts Diffusion Model with a BiMamba–Transformer hybrid backbone and Guidance-Free Training strategy, while the Appearance Expert adopts a decoupled Kinematic–Aesthetic fine-tuning strategy. To better benchmark this task, we curate a large-scale dataset, and design a motion–appearance evaluation protocol. Extensive experiments demonstrate the superiority of MACE-Dance and of its Motion and Appearance Experts. For future work, we plan to extend MACE-Dance with textual descriptions to enable more interactive and flexible dance generation, and improve system-level efficiency to support low-latency authoring and real-time user feedback.
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China under Grants 62436010, 72572090, 62572474 and 62172421, and in part by Tsinghua University School of Economics and Management Research Grant.